Find the intercepts of the given quadratic function.
​f(x) = 5x^2

Answer by ikleyn(52787)

.
y-intercept is at x = 0.

Then f(x) = 0^2 = 0.

So, y-intercept is this point (x,y) = (0,0)   (the origin of the coordinate system).



x-intercept is at f(x) = 0,   or, which is the same,  x^2 = 0,  which gives  x = 0.
Then f(x) = 0^2 = 0.

Thus, y-intercept is this point (x,y) = (0,0)   (the origin of the coordinate system).



In this problem the point of x-intercept is the same as the point of y-intercept.